Need help with questions 7,8,9 &10. Thank you

Now assume a short-run situation where there are a number of different firms using each technique of production, and they are competing within the same industry. The quality of the product is completely standardized; these firms compete only on the basis of how cheaply they can produce it. They can sell output at $93 per unit.

4. Firms using the Original technique of production will find the level of output where profit is maximum (or losses minimum). The difference between revenue and costs, or profit, is $________ (Use negative sign if profit is negative).

            when producing at level of output = ________.

5. Firms using the Challenger technique of production will find the level of output where profit is maximum (or losses minimum). The difference between revenue and costs, or profit, is $_______ (Use negative sign if profits are negative).

   when producing at level of output = _________.

Now consider the long-run situation remembering that positive profits will attract more competitors to the industry and losses will repel firms from the industry.

6. Why will one technique of production prevail over the other? (Identify the prevailing technique and the reason why it prevailed).

7 & 8. How low will price need to fall before firms using the less efficient technique of production will go out of business in the long-run? Why is the threshold price at that particular level?

9 & 10. Ultimately, when all market adjustment is complete, what will be the price in this market?

7 & 8. How low will price need to fall before firms using the less efficient technique of production will go out of business in the long-run? Why is the threshold price at that particular level?

9 & 10. Ultimately, when all market adjustment is complete, what will be the price in this market?

raph 2 output TC TFC TVC 2 2 1 13.64 1023.00-172.00 12.501116.00-269.00 4 1235115.42 190.00102.92 1600 1400 0 12345 6 78 9 10 11 12 -TC_TFC_-TVCATC_MC_AVC_ AFCTRraph 1 output TC TFC TVC ATC MC AVC AFC TR PROFIT 0 100 190 270 100 100 100 100 100 100 100 100 100 100 100 100 0 90 190.00 170 135.00 240 113.33 300 100.00 94.00 91.67 91.43 93.75 97.78 930 103.00 1100109.09 90.00100.00 85.00 80.00 75.00 74.00 75.00 77.14 81.25 86.67 93.00 93.00 50.00186.00 33.33 279.00 25.00372.00 20.00 465.00 16.67 558.00 14.29 651.00 12.50 744.00 97.00 84.00 61.00 28.00 5.00 8.00 11.00 6.00 11.11 837.0043.00 10.00 930.00 -100.00 9.09 1023.00 177.00 2 4 470 370 550 750 880 1030 650 110 780 130 10 170 100.00 1400 1200 1000 0 1 23456 78910 11 -TC_TFC_-TVCATC_MC_AVC_ AFCTR
Answer #1

7 & 8 )

Price will keep falling until it reaches the point at which firm will no longer be able to generate operating profit. This threshold point is known as shut down condition in production theory.

we know,

Revenue = P × Q

Per unit revenue by selling Q amount of goods will be,

rac{P imes Q}{Q} = P

We have ,

Total Cost = Fixed Cost + Variable Cost

Total Avg Cost = Total Cost / Q

Avg Variable Cost = Variable Cost / Q

Once we start producing a good we have already paid the fixed cost ( eg, cost of setting up plant) but it is the variable cost (changes with amount of goods and services produced) which will vary as we change the amount of goods produced.

So, price is the per unit revenue that we generate and AVC is the per unit variable cost that incurs. We will keep producing as long as the per unit profit > 0.

That is P - AVC > 0, which implies P > AVC.

This, P = AVC = MC is known as the shut down condition.

why?

Because generating a operating profit means we are reducing the losses (= fixed cost ) that we are incurring as long as this operating profit will be greater than zero we will produce but not at zero, since then it no longer will compensate for losses incurred due to fixed costs.

8 & 9 )

In the long run all the firms will be producing at the profit = 0, since new firms will keep entering the market as long as there is a scope of positive profit.

This happens when P = MC = AC

such that marginal profit = (P-MC) = 0 for profit maximizing condition and,

Avg Profit = P-AC = 0 for "zero" profit condition.

why?

Because If avg profit = 0,

tota profit = avqpro fit × Q = 0

Firms are still maximizing there output decision even they are making zero profit in long run.

thus prices will keep adjusting until it reaches the point which satisfies the above condition.